### v1

@@ 4 A Structured Neural Network (StNN) for Dynamical Systems
4.1 Structured Neural Network Architecture
@@
We start the section with the forward propagation of the StNN followed by its architecture. The forward propagation of the StNN leverages the factorization equation (5) through specialized layers that incorporate diagonal matrices and recursive operations derived from the factorization of the Hankel operator. This approach is complemented by a layer-by-layer computation process utilizing matrix-vector products, which enables the StNN to achieve enhanced states effectively.
Proposition 4.1.
Let ${\bf x}_{0}\in\mathbb{R}^{m\times 1}$ be the input vector, ${\bf x}_{4}\in\mathbb{R}^{m\times 1}$ be the output vector, and $m$ be the number of nodes in each layer of a neural network. Let the output between the $(i-1)$-th and $i$-th hidden layer be given by:
(10) ${\bf x}_{i}=\sigma_{i}(W_{i,i-1}{\bf x}_{i-1}+{\bf b}_{i})$
where $i:=\{1,2,3,4\},W_{i,i-1}$ is the weight matrix connecting the $(i-1)$-th layer to the $i$-th layer, ${\bf b}$ represents the bias vector, and $\sigma$ is the activation function. Then, we can design a StNN to predict states ${\bf x}_{k+1}$ from ${\bf x}_{k}$ using the weight matrices defined via $W_{1,0}\in\mathbb{R}^{2pm\times m},\quad W_{2,1}\in\mathbb{R}^{pm\times 2pm},\quad W_{3,2}\in\mathbb{R}^{pm\times pm},\quad W_{4,3}\in\mathbb{R}^{m\times pm},$ and their $p$ number of parallel sub-weight matrices, denoted as $w_{i,i-1}$, with the following structured weight matrices
$W_{1,0}=\begin{bmatrix}w_{1,0}\\ w_{1,0}\\ \vdots\\ w_{1,0}\\ \end{bmatrix}_{2pm\times m},W_{2,1}=\begin{bmatrix}w_{2,1}&0&...&0\\ 0&w_{2,1}&...&0\\ \vdots&\vdots&&\vdots\\ 0&0&...&w_{2,1}\end{bmatrix}_{pm\times 2pm},$
$W_{3,2}=\begin{bmatrix}w_{3,2}&0&...&0\\ 0&w_{3,2}&...&0\\ \vdots&\vdots&&\vdots\\ 0&0&...&w_{3,2}\end{bmatrix}_{pm\times pm},W_{4,3}=\begin{bmatrix}w_{4,3}&w_{4,3}&...&w_{4,3}\end{bmatrix}_{m\times pm},$
where $w_{1,0}=\mathfrak{F}^{*}_{2m}[J]_{2m\times m}\hat{D}_{m}\in\mathbb{R}^{2m\times m}$, $w_{2,1}=[J^{T}]_{m\times 2m}\mathfrak{F}_{2m}\breve{{\bf D}}_{2m}\in\mathbb{R}^{m\times 2m}$,$w_{3,2}=\tilde{\bf I}_{m}\hat{\bf D}_{m}\in\mathbb{R}^{m\times m}$, and $w_{4,3}=D_{m}\in\mathbb{R}^{m\times m}$.
Proof 4.2.
@@
Let us define the sub-matrices $w_{1,0}\in\mathbb{R}^{2m\times m}$, $w_{2,1}\in\mathbb{R}^{m\times 2m}$,$w_{3,2}\in\mathbb{R}^{m\times m}$, and $w_{4,3}\in\mathbb{R}^{m\times m}$ based on the factorization of the Hankel operator (5) to design layers and learn weights for the proposed network. We begin by grouping the matrices in the factorization (5) into three distinct groups, ensuring each group corresponds to the weight matrices connecting $(i-1)$-th layer to the $i$-th layer. In the first hidden layer, we define $j$ parallel sub-weight matrices s.t. $w_{1,0}=\mathfrak{F}^{*}_{2m}[J]_{2m\times m}\hat{D}_{m}$, to learn the weight matrix $W_{1,0}$. Next, the sub-weight matrices connecting the first and second hidden layers are defined by $w_{2,1}=[J^{T}]_{m\times 2m}\mathfrak{F}_{2m}\breve{{\bf D}}_{2m}$ which are utilized to learn the weight matrix $W_{2,1}$. Next, the sub-weight matrices between the second and third hidden layers are defined as $w_{3,2}=\tilde{\bf I}_{m}\hat{\bf D}_{m}$, and we utilize those to learn the weight matrix $W_{3,2}$. After the third hidden layer, the sub-weight matrices connecting the last hidden layer to the output layer are represented as diagonal matrices, i.e., $w_{4,3}=D_{m}$. Consequently, a linear transformation is applied to combine the outputs of the sub-weight matrices to learn the weight matrix $W_{4,3}$. In addition to these weight matrices, we have frozen all identity and zero matrices in the factorization of the Hankel operator (5) at each network layer. This will enable us to develop a lightweight model. Also, we have not shared or reused matrices among different layers, ensuring that no additional matrices contribute to the network architecture. With this configuration of parallel sub-weight matrices and frozen matrices, along with the propagation described in equation (10), we efficiently train the weight matrices of the StNN using a lightweight model.
@@
To illustrate the advantages of our proposed network architecture over the feed-forward neural network (FFNN), we will present the structure of the StNN alongside the FFNN followed by the flop count, as detailed in Table 1.
Table 1: The StNN and FFNN architectures are designed for the layer-wise comparison of weight matrices, biases, total number of parameters, and flop counts. The value $p$ denotes the number of parallel sub-weight matrices designed for the values $p=1,2,4,6$, which correspond to four distinct StNN models. These sub-matrices are used to construct and learn weight matrices $W_{i,i-1}$ that connect the $(i-1)$-th layer to the $i$-th layer for $i=1,2,3,4$.
Weight Sub Number of Weights Biases Total flop
Matrix Weight Parallel Sub Parameters Count
Matrices Weight Matrices
StNN (Structured Neural Network)
$W_{1,0}$
$[\mathbf{\hat{D}}]_{4\times 4}$
$[\mathfrak{F_{2}}]_{2\times 2}$
$[\mathbf{H]_{8\times 8}}$
$[\mathbf{H]_{4\times 4}}$
$p$
$2p$
$p$
$p$
$4p$
$8p$
$4p$
$2p$
$8p$ $26p$ $68p$
Total - $\mathbf{18p}$ $\mathbf{8p}$ $\mathbf{26p}$ $\mathbf{68p}$
$W_{2,1}$
$[\mathbf{\hat{D}}]_{8\times 8}$
$[\mathfrak{F_{2}}]_{2\times 2}$
$[\mathbf{H]_{8\times 8}}$
$[\mathbf{H]_{4\times 4}}$
$p$
$2p$
$p$
$p$
$8p$
$8p$
$4p$
$2p$
$4p$ $26p$ $68p$
Total - $\mathbf{22p}$ $\mathbf{4p}$ $\mathbf{26p}$ $\mathbf{68p}$
$W_{3,2}$
$[\mathbf{\hat{D}}]_{4\times 4}$
$p$
$4p$
$4p$ $8p$ $8p$
$W_{4,3}$
$[\mathbf{\hat{D}}]_{4\times 4}$
$p$
$4p$
$4p$ $8p$ $412p-$ 4
Total - - $\mathbf{48p}$ $\mathbf{20p}$ $\mathbf{68p}$ $\mathbf{156p-4}$
FFNN (Feed-forward Neural Network)
$W_{1,0}$ $[W]_{30\times 3}$ $90$ $30$ $120$ $180$
$W_{2,1}$ $[W]_{30\times 30}$ $900$ $30$ $930$ $1800$
$W_{3,2}$ $[W]_{30\times 30}$ $900$ $30$ $930$ $1800$
$W_{4,3}$ $[W]_{3\times 30}$ $90$ $3$ $93$ $180$
Total - - $1980$ $93$ $2073$ $3960$
@@
4.2 Structured Neural Network Approach to Predict Trajectories of Dynamical Systems
@@
(11) $\frac{dx}{dt}=\sigma(y-x),$ $\frac{dy}{dt}=x(\rho-z)-y,$ $\frac{dz}{dt}=xy-\beta z$
@@
To obtain the evolution of the Lorenz system, we generate a wide range of initial conditions, denoted by vector ${\bf x}_{0}$, and track the trajectories over time. We advance the initial conditions with a sampling time interval of $\Delta t$, which is not the actual time step. The next step is to acquire the matrices that represent the inputs and outputs of the system at states ${\bf x}_{k}$ and ${\bf x}_{k+1}$, respectively, with sample increments of $\Delta t$, which are correlated to ${\bf X}$ and ${\bf{X^{\prime}}}$, respectively. These matrices are obtained by utilizing the trajectories that have been trained over time through the learned Hankel operator ${\bf H}$. Thus to capture the evolution of the non-linear nature of the dynamical systems, we use StNN and FFNN with 5 layers (when including the input, output, and 3 hidden layers) and different number of nodes in each layer while imposing the structure to the network using Propositions 2.5 and 4.1, in order to carry on the forward propagation. The network will be trained on trajectories based on (9) to predict states in future time for any given initial conditions. The network will utilize activation functions such as Tanh and Sigmoid in the first two hidden layers and ReLU as the activation function of the third hidden layer to incorporate the dynamics of the system. As a result, we derive a new set of spatiotemporal data to generate future predictions from ${\bf x}_{k}$ to ${\bf x}_{k+1}$.
Additionally, we evaluate the training performance over $e$ epochs, using the loss function based on Proposition 3.1 s.t.
$L({\bf x}_{k},{\bf x}_{k+1}):=\|{\bf{X^{\prime}}}-\widehat{\bf H}{\bf X}\|^{2},$
and hence validate the trajectory data of the trained model against the dynamical model using the best-fit time advanced state-based Hankel operator $\widehat{\bf H}$.
5 Numerical Simulations: Learn, Update, and Predict States
@@
5.1 Numerical Simulations: Lotka-Volterra Model to Learn and Predict Dynamics
In this section, we show numerical simulations to determine the time evolution of the Lotka-Volterra model for different environments, where each environment is described by a set of system parameters $\alpha,\beta,\gamma$, and $\delta$. In this experiment, we also draw comparisons with recently proposed model for dynamical systems, LEADS [49]. LEADS is a framework that leverages the commonalities and discrepancies among known environments to improve model generalization, using separate model components that focus either on global or environmental-specific dynamics. Following the experimental setup from LEADS, we consider 10 possible environments and generate trajectories each with 20 data points in time, $t_{k}=0.0,0.5,1.0,\dots,9.5$. For training, we sample 8 trajectories from each environment. Each environment has a unique set of system parameters, while each trajectory has a unique set of initial conditions. At evaluation, the models are tested on 32 trajectories from each environment. The model receives $x{(t_{k})}$, $y(t_{k})$, $t_{k}$ and the environment passed as a unique integer which parametrizes the system parameters. The goal is to predict $x{(t_{k+1})},y{(t_{k+1})}$ as outputs of the model. During the evaluation, the model only receives the initial conditions $x(t_{0}),y(t_{0}),t_{0}=0$, and environment specifier, performing an autoregressive rollout to predict all future $x,y$. The results of this experiment are summarized in Table 2 and examples of predictions are provided in Figure 2.
Table 2: Test results of the StNN on the Lotka-Volterra equations. Baseline experiments with the LEADS model [49] show that the proposed approach is able to obtain remarkable accuracy with very few parameters.
@@ 5.1 Numerical Simulations: Lotka-Volterra Model to Learn and Predict Dynamics
Although LEADS was designed with novel elements to improve generalization across environments, we observe that large deviations from the truth may arise in some instances, illustrated in Figure 2 (b). Meanwhile, StNN predicts dynamics which remain close to the ground truth. Additionally, Table 2 illustrates several advantages of the StNN in parameter complexity and training time requirements. While LEADS has nearly 100,000 parameters, StNN is able to achieve a competitive error with a remarkable 388 trainable parameters. As a result, StNN may also be trained an entire order of magnitude faster than the competing approach. This experiment underlines the advantages of structured matrices with learnable parameters, as we propose in this work.
5.2 Numerical Setup for the Chaotic Lorenz System
@@
In this section, we analyze the performance of the StNN architecture compared to FFNN, SINDy, and HAVOK using the chaotic Lorenz system. To conduct these simulations, the Lorenz system, characterized by the differential equations (11), was used to produce time-series data based on the parameters $\sigma=10$, $\beta=\frac{8}{3}$, and $\rho=28$ with a time step of $dt=0.01$ at the duration of $T=8$, i.e. each trajectory consists of 800 data points. Furthermore, we obtained 100 such trajectories by perturbing the nominal initial state, i.e, $[x(t_{0}),y(t_{0}),z(t_{0})]=[0,1,20]$ with random uniform noise of magnitude 1. The odeint function from the SciPy library was used to numerically integrate each perturbed trajectory, guaranteeing high accuracy with relative and absolute tolerances set to $1\times 10^{-12}$. Figure 3 shows the trajectories of the Lorenz attractor, along with 100 randomly perturbed trajectories originating from the initial states, represented by red dots.
Figure 3: Visualization of Lorenz system trajectories with random perturbations in initial conditions. The red dots represent the perturbed starting points, and the lines illustrate the chaotic evolution of the system.
As shown in Figure 4, we observed that the trained StNN reproduced the chaotic dynamics of the system with high accuracy and reduced complexity as compared to the FFNN. This is evident from the training and validation MSE comparison between the StNN and the FFNN in Figure 4. The SNN maintains a more resource-efficient architecture while demonstrating a smoother training MSE curve, which shows steady performance gains over time. This demonstrates how the chosen StNN can balance computational efficiency and accuracy, making it a competitive alternative to the FFNN. However, the FFNN, having a larger number of trainable weights, demonstrates greater flexibility and achieves lower MSE rates. The reason is that the increased capacity of the FFNN allows it to better approximate the target function. Conversely, StNN’s reduced parameterization naturally imposes constraints on the network’s flexibility, resulting in a slightly higher overall MSE. Despite this, the StNN’s lightweight architecture offers significant advantages in terms of computational efficiency, making it an appealing choice for scenarios where inference speed and resource constraints are critical considerations.
(a) FFNN and StNN training MSE over steps
(b) FFNN and StNN validation MSE over steps
Figure 4: Training and validation MSE comparison for FFNN and StNN models over states evaluated at time steps. Each step represents a randomly sampled mini-batch from the dataset, with 640 steps:= 1 training epoch. The plots illustrate the convergence behavior, showing how the MSE evolves as training progresses. The plot (a) depicts training MSE, while the plot (b) shows validation MSE. While the FFNN (blue) shows a lower minimum MSE than the StNN (red) during both training and validation, The dotted lines represent the minimum recorded MSE for each model. Moreover, Table 3 illustrates that the StNN is more lightweight than the FFNN, resulting in reduced complexity.
5.3 Numerical Simulations of NNs: Learn and Update
@@
The StNN was implemented using a feedforward architecture with input, output, and three hidden layers, as explained in Section 4.2. This model effectively combines activation functions ($\tanh$, $Sigmoid$, and $ReLU$) to capture the complex non-linear dynamics of the Lorenz system. The input and output dimensions were set to 4 by padding the state variables ($x,y,z$) with 0, i.e., ($x,y,z,0$)to match the dimensions.
@@
An 80,000-sized randomly generated dataset was divided into input-output pairs, with each input being a state vector $[x_{1}(t_{k}),x_{2}(t_{k}),x_{3}(t_{k}),0]$ and the output being the subsequent state vector $[x_{1}(t_{k+1}),x_{2}(t_{k+1}),x_{3}(t_{k+1}),0]$. The input and output datasets were mini-batched $1000$ to ensure efficient mini-batch-wise training[21] We split the dataset into training and validation, i.e., 80% of the dataset was allocated for training, while the remaining 20% was reserved for validation. To enhance readers’ understanding of the theoretical foundation and its connection to the StNN learning algorithm, we direct readers to the code StNN-Dynamical-Systems.
We utilized the Levenberg-Marquardt algorithm implemented in PyTorch by Di Marco [31]. This implementation enables efficient optimization for training neural networks by combining the advantages of gradient descent and Newton’s method. The training process was conducted over 20 epochs with 640 steps for each epoch, utilizing the high convergence rate of the Levenberg-Marquardt method for non-linear regression tasks. First, we present the convergence performance of the StNN in comparison with the FFNN while varying the number of sub-weight matrices $p$. As discussed in Section 4, the StNN can be represented using multiple parallel sub weight matrices, denoted as $p$. The selection of an appropriate value for $p$ should be guided by the desired MSE and the specific complexity requirements of the system. A summary of the training and validation performance for various StNN models with different $p$ values is provided in Table 3. The flop savings percentages are calculated in comparison to the FFNN using
@@
Table 3: We show the training and validation performance of StNN for different numbers of parallel $p$ sub-weight matrix configurations. This table summarizes the impact of varying $p$ on Mean Squared Error (MSE), model weights, and computational complexity (flops). Savings percentages are calculated in comparison to the FFNN using equation (12).
p Final Parameter Saving flop Saving
Training MSE compared to FFNN in compared to FFNN in
Table 1 Table 1
1 $1.49\times 10^{-4}$ $96\%$ $96\%$
2 $2.47\times 10^{-4}$ $94\%$ $92\%$
4 $2.67\times 10^{-6}$ $87\%$ $84\%$
6 $5.26\times 10^{-6}$ $81\%$ $76\%$
Smaller $p$ values lead to higher MSEs, reflecting the trade-off between model simplicity and accuracy. While smaller $p$ values reduce the number of weights and floating-point operations, they also result in less precise predictions over time. Conversely, larger $p$ values yield significantly lower MSEs, highlighting their excellent predictive accuracy. However, this improvement comes at the cost of increased computational complexity, as seen in the greater number of weights and flops required. Interestingly, when compared to other StNN configurations, the best-performing StNN model is $p=6$, which achieves a significant reduction in the error on the dataset. Furthermore, compared to the FFNN, this StNN achieves a reduction of roughly $81\%$ in the number of weights and $76\%$ in floating-point operations, demonstrating a significant advantage in computational and parameter complexity.
5.4 Numerical Simulations of NNs: Time Evolution
Once the StNN and FFNN are trained and updated on the trajectory data in Section 5.3, the non-linear dynamical model describing the Lorenz system could be to map the states from ${\bf x}_{k}$ to ${\bf x}_{k+1}$ and hence to predict the future states from an initial state.
(a) FFNN autoregressive prediction over 600-time steps
(b) StNN autoregressive prediction over 600-time steps

### v2

@@ III A Structured Neural Network (StNN) for Dynamical Systems
III-A Structured Neural Network Architecture
We start the section with the forward propagation of the StNN, followed by its architecture. The forward propagation of the StNN leverages the revised factorization equation (8), i.e.,
$\widehat{\bf H}\approx\tilde{\bf I}_{n}[{\bf J}^{T}]_{n\times r}{F}_{r}\breve{\bf D}_{r}{F}_{r}[{\bf J}]_{r\times n},$ (9)
@@ III-A Structured Neural Network Architecture
where symmetric sub-weight matrices $F_{r}\in\mathbb{R}^{r\times r}$ approximate ${\bf\mathfrak{F}}_{r}$, through customized layers that incorporate diagonal matrices and recursive FFT-like factorization. In other words, the layers of the network are designed to execute with real-valued inputs based on the factorization equation (8) followed by a divide-and-conquer technique of $F_{r}$ as a real-valued replacement for ${\bf\mathfrak{F}}_{r}$ through structured learning. This approach is complemented by a layer-by-layer computation process utilizing matrix-vector products, which enables the StNN to achieve enhanced states effectively.
Proposition 4.
[…]
${\bf x}_{i}=\sigma_{i}(W_{i,i-1}{\bf x}_{i-1}+{\bf b}_{i})$ (10)
[…]
$W_{1,0}=\begin{bmatrix}w_{1,0}\\ w_{1,0}\\ \vdots\\ w_{1,0}\\ \end{bmatrix}_{2pn\times n},W_{2,1}=\begin{bmatrix}w_{2,1}&0&...&0\\ 0&w_{2,1}&...&0\\ \vdots&\vdots&&\vdots\\ 0&0&...&w_{2,1}\end{bmatrix}_{pn\times 2pn},$
$W_{3,2}=\begin{bmatrix}w_{3,2}&0&...&0\\ 0&w_{3,2}&...&0\\ \vdots&\vdots&&\vdots\\ 0&0&...&w_{3,2}\end{bmatrix}_{pn\times pn},$
$W_{4,3}=\begin{bmatrix}w_{4,3}&w_{4,3}&...&w_{4,3}\end{bmatrix}_{n\times pn},$
where $w_{1,0}={F}_{r}[{\bf J}]_{r\times n}\in\mathbb{R}^{r\times n}$, $w_{2,1}=[{\bf J}^{T}]_{n\times r}{F}_{r}\hat{D}_{r}\in\mathbb{R}^{n\times r}$,$w_{3,2}=\tilde{\bf I}_{n}\in\mathbb{R}^{n\times n}$, $w_{4,3}=D_{n}\in\mathbb{R}^{n\times n}$, ${F}_{r}=P_{r}^{T}\begin{bmatrix}{F}_{n}&\\ &{F}_{n}\end{bmatrix}H_{r}$ with random initialization of $\times 2$ 2 weight matrices, $P_{r}$ is an even-odd permutation matrix, ${H}_{r}=\left[\begin{array}[]{rr}{\bf I}_{n}&{\bf I}_{n}\\ \grave{D}_{n}&-\grave{D}_{n}\\ \end{array}\right]$, $r=2n$, and $D_{n}$, $\hat{D}_{n}$ and $\grave{D}_{n}\in\mathbb{R}^{n\times n}$ are randomized diagonal weight matrices.
Proof.
@@
Let us define the sub-matrices $w_{1,0}\in\mathbb{R}^{r\times n}$, $w_{2,1}\in\mathbb{R}^{n\times r}$,$w_{3,2}\in\mathbb{R}^{n\times n}$, and $w_{4,3}\in\mathbb{R}^{n\times n}$ based on the factorization of the Hankel operator (8) followed by (9) and a divide-and-conquer technique to design layers and learn weights for the proposed network. We begin by grouping the matrices in the factorization (8) followed by (9) into three distinct groups, followed by a random diagonal weight matrix, ensuring each group corresponds to the weight matrices connecting $(i-1)$-th layer to the $i$-th layer. In the first hidden layer, we define $j$ parallel sub-weight matrices s.t. $w_{1,0}=F_{r}[{\bf J}]_{r\times n}$, to learn the weight matrix $W_{1,0}$. Next, the sub-weight matrices connecting the first and second hidden layers are defined by $w_{2,1}=[{\bf J}^{T}]_{n\times r}F_{r}\breve{\bf D}_{r}$ which are utilized to learn the weight matrix $W_{2,1}$. Next, the sub-weight matrices between the second and third hidden layers are defined as $w_{3,2}=\tilde{\bf I}_{n}$, and we utilize those to learn the weight matrix $W_{3,2}$. After the third hidden layer, the sub-weight matrices connecting the last hidden layer to the output layer are represented as diagonal matrices, i.e., $w_{4,3}=D_{n}$. Consequently, a linear transformation based on diagonal weight matrices is applied to combine the outputs of the sub-weight matrices to learn the weight matrix $W_{4,3}$. In addition to these weight matrices, we have frozen all identity and zero matrices in the factorization of the Hankel operator (8) followed by (9) at each network layer. This will enable us to develop a lightweight model. Also, we have not shared or reused matrices among different layers, ensuring that no additional matrices contribute to the network architecture. With this configuration of parallel sub-weight matrices and frozen matrices, along with the propagation described in equation (10), we efficiently train the weight matrices of the StNN using a lightweight model. ∎
@@
To illustrate the advantages of our proposed network architecture over the feed-forward neural network (FFNN), we will present the structure of the StNN alongside the FFNN followed by the flops count, as detailed in Table V.
@@
III-B Structured Neural Network Approach to Predict Trajectories of Dynamical Systems
@@
$\frac{dx}{dt}=\sigma(y-x),$ $\frac{dy}{dt}=x(\rho-z)-y,$ $\frac{dz}{dt}=xy-\beta z$ (11)
@@
To obtain the evolution of the Lorenz system, we generate a wide range of initial conditions, denoted by vector ${\bf x}_{0}$, and track the trajectories over time. We advance the initial conditions with a sampling time interval of $\Delta t$, which is not the actual time step. The next step is to acquire the matrices that represent the inputs and outputs of the system at states ${\bf x}_{k}$ and ${\bf x}_{k+1}$, respectively, with sample increments of $\Delta t$, which are correlated to ${\bf X}$ and ${\bf{X^{\prime}}}$, respectively. These matrices are obtained by utilizing the trajectories that have been trained over time through the learned Hankel operator $\widehat{\bf H}$. Thus to capture the evolution of the non-linear nature of the dynamical systems, we use StNN and FFNN with 5 layers (when including the input, output, and 3 hidden layers) and different number of nodes in each layer while imposing the structure to the network using Propositions 3 and 4, in order to carry on the forward propagation. The network will be trained on trajectories based on (7) to predict states in future time for any given initial conditions. The network will utilize activation functions such as Tanh and Sigmoid in the first two hidden layers and ReLU as the activation function of the third hidden layer to incorporate the dynamics of the system. As a result, we derive a new set of spatiotemporal data to generate future predictions from ${\bf x}_{k}$ to ${\bf x}_{k+1}$.
Additionally, we evaluate the training performance over $e$ epochs, using the loss function based on Proposition 1 s.t.
$L({\bf x}_{k},{\bf x}_{k+1}):=\frac{1}{m_{b}\times n}\|{\bf{X^{\prime}}}-\widehat{\bf H}{\bf X}\|^{2}_{F}+\sum_{l=1}^{4}\alpha_{l}\|W_{l,l-1}\|_{\eta},$ (12)
where $m_{b}$ represents the mini-batch size and $\alpha_{l}$ denotes the regularization hyperparameter that must be tuned for each layer and hence validate the trajectory data of the trained model against the dynamical model using the best-fit time-advanced state-based Hankel operator $\widehat{\bf H}$.
IV Numerical Simulations: Learn, Update, and Predict States
@@
IV-A Numerical Simulations: Lotka-Volterra Model to Learn and Predict Dynamics
In this section, we show numerical simulations to determine the time evolution of the Lotka-Volterra model for different environments, where each environment is described by a set of system parameters $\alpha,\beta,\gamma$, and $\delta$. In this experiment, we also draw comparisons with recently proposed model for dynamical systems, LEADS [44]. LEADS is a framework that leverages the commonalities and discrepancies among known environments to improve model generalization, using separate model components that focus either on global or environmental-specific dynamics. Following the experimental setup from LEADS, we consider 10 possible environments and generate trajectories each with 20 data points in time, $t_{k}=0.0,0.5,1.0,\dots,9.5$. For training, we sample 8 trajectories from each environment. Each environment has a unique set of system parameters, while each trajectory has a unique set of initial conditions. At evaluation, the models are tested on 32 trajectories from each environment. The model receives $x{(t_{k})}$, $y(t_{k})$, $t_{k}$ and the environment passed as a unique integer which parametrizes the system parameters. The goal is to predict $x{(t_{k+1})},y{(t_{k+1})}$ as outputs of the model. During the evaluation, the model only receives the initial conditions $x(t_{0}),y(t_{0}),t_{0}=0$, and environment specifier, performing an autoregressive rollout to predict all future $x,y$. The results of this experiment are summarized in Table I and examples of predictions are provided in Figure 2.
TABLE I: Test results of the StNN on the Lotka-Volterra equations. Baseline experiments with the LEADS model [44] show that the proposed approach is able to obtain remarkable accuracy with very few parameters.
@@ IV-A Numerical Simulations: Lotka-Volterra Model to Learn and Predict Dynamics
Although LEADS was designed with novel elements to improve generalization across environments, we observe that large deviations from the truth may arise in some instances, illustrated in Figure 2 (b). Meanwhile, StNN predicts dynamics which remain close to the ground truth. Additionally, Table I illustrates several advantages of the StNN in parameter complexity and training time requirements. While LEADS has nearly 100,000 parameters, StNN is able to achieve a competitive error with a remarkable 388 trainable parameters. As a result, StNN may also be trained an entire order of magnitude faster than the competing approach. This experiment underlines the advantages of structured matrices with learnable parameters, as we propose in this work.
IV-B Numerical Setup for the Chaotic Lorenz System
@@
In this section, we analyze the performance of the StNN architecture compared to FFNN, SINDy, and HAVOK using the chaotic Lorenz system. To conduct these simulations, the Lorenz system, characterized by the differential equations (11), was used to produce time-series data based on the parameters $\sigma=10$, $\beta=\frac{8}{3}$, and $\rho=28$ with a time step of $dt=0.01$ at the duration of $T=8$, i.e. each trajectory consists of 800 data points. Furthermore, we obtained 100 such trajectories by perturbing the nominal initial state, i.e, $[x(t_{0}),y(t_{0}),z(t_{0})]=[0,1,20]$ with random uniform noise of magnitude 1. The odeint function from the SciPy library was used to numerically integrate each perturbed trajectory, guaranteeing high accuracy with relative and absolute tolerances set to $1\times 10^{-12}$.
@@
The StNN was implemented using a feedforward architecture with input, output, and three hidden layers, as explained in Section III-B. This model effectively combines activation functions ($Tanh$, Leaky-ReLU, and $ReLU$) to capture the complex non-linear dynamics of the Lorenz system. The input and output dimensions were set to 4 by padding the state variables ($x,y,z$) with 0, i.e., ($x,y,z,0$)to match the dimensions.
@@
An 80,000-sized randomly generated dataset was divided into input-output pairs, with each input being a state vector $[x_{1}(t_{k}),x_{2}(t_{k}),x_{3}(t_{k}),0]$ and the output being the subsequent state vector $[x_{1}(t_{k+1}),x_{2}(t_{k+1}),x_{3}(t_{k+1}),0]$. The input and output datasets were mini-batched 100010001000 to ensure efficient mini-batch-wise training[45]. We split the dataset into training and validation, i.e., 80% of the dataset was allocated for training, while the remaining 20% was reserved for validation. We utilized the Levenberg-Marquardt algorithm implemented in PyTorch by Di Marco [46]. This implementation enables efficient optimization for training neural networks by combining the advantages of gradient descent and Newton’s method. The training process was conducted over 20 epochs with 640 steps for each epoch, utilizing the high convergence rate of the Levenberg-Marquardt method for non-linear regression tasks. We set the hyperparameter value $\alpha_{2}=1\times 10^{-7}$ and $\alpha_{1},\alpha_{3},\alpha_{4}=0$. The main reason for this selection is that we want to minimize the loss function (12) with the order of $10^{-6}$. Therefore, as the error approaches this order, we aim to balance the error term with the nuclear norm regularization term to enforce a low-rank structure in the StNN through the loss function 12. To enhance readers’ understanding of the theoretical foundation and its connection to the StNN learning algorithm, we direct readers to the code StNN-Dynamical-Systems.
A summary of the training and validation performance for various StNN models with different $p$ values is provided in Table II. The flops or parameter savings percentages are calculated in comparison to the FFNN using
@@
#$\frac{\#\text{FFNN(flops/parameters)}-\#\text{StNN(flops/parameters)}}{\#\text{FFNN(flops or parameters)}}\times 100\%.$
TABLE II: We show the training and validation performance of StNN for different numbers of parallel $p$ sub-weight matrix configurations based on Table V. This table summarizes the impact of varying $p$ based on the loss function (12) taken as training error, model weights, and computational complexity (flops). Savings percentages are calculated in the comparison of StNN to the FFNN using equation (IV-B).
p Training Training Parameter Saving flops Saving
Error (12) Time StNN vs FFNN StNN vs FFNN
[…]
Once the StNN and FFNN are trained and updated on the trajectory data, the non-linear dynamical model describing the Lorenz system could be to map the states from ${\bf x}_{k}$ to ${\bf x}_{k+1}$ and hence to predict the future states from an initial state.
(a) FFNN autoregressive prediction over 5000-time steps
(b) StNN autoregressive prediction over 500-time steps